OpenStudy (anonymous):

Totally lost on these three.... "Suppose f(x) = -3x + 2 and g(x) = 1/2x - 1. Find each value" 17) 3g(4) 18) g(-2)/f(3) 19) f(-1)/g(5)

OpenStudy (helder_edwin):

u have \[ \large g(x)=\frac{1}{2x-1} \] to find g(5) replace x=5 in every x to right \[ \large g(5)=\frac{1}{2\cdot5-1}=\frac{1}{10-1}=\frac{1}{9} \]

OpenStudy (helder_edwin):

does this help?

OpenStudy (anonymous):

It's actually g(x) = 1/2 times x, minus one. Not one over 2x-1

OpenStudy (helder_edwin):

it doesn't matter. take it as an example. do u understand the substitution?

OpenStudy (helder_edwin):

if \[ \large g(x)=\frac{1}{2}x-1 \] then if x=7 \[ \large g(7)=\frac{1}{2}\cdot7-1=\frac{7}{2}-1=\frac{5}{2} \]

OpenStudy (anonymous):

Yeah, I get that. The trouble is when they add the fractions in on 18 and 19. I'm just totally lost....

OpenStudy (helder_edwin):

\[ \large \frac{g(-2)}{f(3)}=\frac{\frac{1}{2}(-2)-1}{-3(3)+2}= \]

OpenStudy (helder_edwin):

it is quite the same. just do it separatedly